1 Overview

Figure 1: Time evolution of the GOES X-ray fluxes. Here we focus on the October 28, 2003 X17 flare. In the bottom plot, the timings of the 5 successive CDS rasters are indicated. CDS observed the two precursor M-class flares F1 and F2 with the third raster.

2 The two pre-events, M-class flares

SOHO CDS observed with the 3rd raster during 09:37-11:02 UT. CDS observed two distinct M-class flares, with strong Fe XIX emission (saturated in the two cores). The first one (F1) was observed during 10:00-10:19 UT, while the second (F2) during 10:33-10:47 UT. The first flare was also observed by TRACE.

3 The X17 flare

Post-flare arcade

Here we show results concerning the 5th CDS observation (12:28-12:52 UT). During this time, TRACE 195 and 1600 Å images show that the 2-ribbon structure and the arcade of loops is approximately stationary, with some apparent flows along the loops.

Figure 2: Sequence of TRACE 195 Å images, showing the two M-class flares. Between 10:24 and 10:49 UT TRACE did not observe. Note the contours of Fe XIX emission of the two flares observed by CDS.

Figure 3: Sequence of TRACE 1600 Å images, showing the two M-class flares. Between 10:24 and 10:49 UT TRACE did not observe.

Figure 4: TRACE 195,1600 Å with contours from SOHO/MDI magnetogram. The M-class flare occurred in the region of mixed polarity. Note also the filament eruption associated with this flare.

Figure 5: SOHO/MDI magnetogram.

Figure 6: CDS monochromatic images at transition region (O V), coronal (Mg X) and flare (Fe XIX) temperatures, during the long CDS scan which observed the two F1,F2 M-class flares. Overlaid are contours of TRACE 1600 Å. Notice the complex structure of the second flare (left) in the enhanced Fe XIX image.

Figure 7: Sequence of TRACE 195 Å images during the X17 flare. The first three images are shown with a different intensity scale, compared to the following ones.

Figure 8: Top: TRACE 195 Å with contours of CIV. Bottom: TRACE 1600 Å Figure 9: Top: SOHO/MDI magnetogram with contours of CIV (showing the ribbons). Bottom: TRACE 195 Å with contours of Fe XIX intensity from CDS. The contours show the hottest regions, where most TRACE emission is dominated by Fe XXIV.

Figure 10: CDS intensity maps with contours of CIV (showing the ribbons). The post-flare loops show emission at all temperatures Figure 11: CDS velocity map in the transition region (O V) with contours of CIV (showing the ribbons). Strong downflows (v ≃ 50 − 100 km/s ) in the ribbons are observed.

Figure 12: CDS velocity maps in coronal lines with contours of C IV (showing the ribbons). Strong downflows (v ≃ 50 − 100 km/s ) in the ribbons are observed.

• The X17 flare presented great complexity. We have used the CDS to understand the temperature structure of the various events.

• Two precursor M-class events have been observed with CDS. The first one was also observed by TRACE, while the second, more important one, was not observed by TRACE. The second one had a complex structure of hot loops.

• During the precursor events the filament becomes activated.

• During the X17 flare the ribbons quickly separate and an arcade system of loops forms. These loops are multi-thermal, with the hottest part in the central region.

• Strong downflows (v ≃ 50 − 100 km/s ), in particular at transition region temperatures are observed in the ribbons.